OpenStudy (anonymous):

Show (sin x+cos x)^2<2

OpenStudy (primeralph):

Check the maxima of the function.

OpenStudy (anonymous):

actually, (cos^2(x) + sin^2(x))^2 = 2 when x = pi/4 perhaps you meant (cos^2(x) + sin^2(x)) <= 2 expanding gives sin^2x + cos^2x + 2cosx sinx using trig identities, this equals to 1 + sin(2x) we know that -1 <= sin(2x) <= 1 0 <= 1 + sin(2x) <= 2

OpenStudy (anonymous):

I meant (cosx + sinx)^2 = 2 when x = pi/4

OpenStudy (anonymous):

No, actually it's |sin x+cos x|< √ ̅ 2

OpenStudy (anonymous):

well when x = pi/4, then the expression equals sqrt(2)

OpenStudy (anonymous):

and what about |a*sin x+b*cos x|<√ ̅ (a^2+b^2)

OpenStudy (anonymous):

what if a=b=0 ?

OpenStudy (anonymous):

I think |a*sin x+b*cos x|<=√ ̅ (a^2+b^2)

OpenStudy (anonymous):

but it doesn't matter because |sin x+cos x|< √ ̅ 2 is false anyway. but |sin x+cos x|<= √ ̅ 2 is true

OpenStudy (anonymous):

yes. it comes <= .I forgot to put it

OpenStudy (anonymous):

actually if a=b=0 then a*sinx and b*cos x is simpliying so that's not the answer.. anyway

OpenStudy (anonymous):

so you want to show that |sin x+cos x|<= √ ̅ 2 ?

OpenStudy (anonymous):

that the first.. then i have to show |a*sin x+b*cos x|<=√ ̅ (a^2+b^2)

OpenStudy (raden):

just hint : a*sin x+b*cos x can be converted to k cos(x-θ), with k = sqrt(a^2 + b^2)

OpenStudy (anonymous):

I give up :|

OpenStudy (raden):

the 1st and the 2nd problem are exactly the samethink. see how to change sin x+cos x into kcos(x-θ) here a = 1 and b = 1, so k = sqrt(1^1+1^2) = sqrt(2) and tanθ = a/b = 1/1 = 1 (here θ in the 1st quadrant) get θ = 45 degrees therefore, |sin x+cos x|<= √ ̅ 2 |sqrt(2) cos(x-θ)} <= √ ̅ 2 and we knowed that -1 <= cos(x-θ) <= 1 and absolut sign makes everthink be positive and zero. so, obvious the LHS would be less than or equal the RHS